OpenStudy (anonymous):

Medal and fan Which cation from the table has the greatest solubility ? BaSO4 Ksp = 1.1*10^-10 CaCO3 Ksp = 2.8*10^-9 CaF2 Ksp = 5.3*10^-9 CuCl Ksp = 1.2*10^-6 I had answered with CuCl, it marked me wrong. Which one would it be then? Thanks(:

OpenStudy (anonymous):

@sweetburger

OpenStudy (somy):

where are cations in the table

OpenStudy (anonymous):

The Ba, Cu, Ca.

OpenStudy (somy):

is it BaSO4 Ksp = 1.1*10^-10 ?

OpenStudy (anonymous):

How? That just lowers the ksp, which makes it less soluble?

OpenStudy (somy):

nah but is it right?

OpenStudy (somy):

i didn't even think of Ksp

OpenStudy (anonymous):

I won't know till the end haha

OpenStudy (somy):

well i do remember learning that solubility of compounds of group 2 increases down the group

OpenStudy (somy):

the one that is on the very bottom is Ba from give cations thus i chose A

OpenStudy (anonymous):

But BaSO4 has the smallest ksp..

OpenStudy (somy):

hmmm don't know really this is actually what i remember

OpenStudy (anonymous):

I just think it has more to do with the points in this lesson rather than previous ones, otherwise they wouldn't give ksp.

OpenStudy (somy):

ik thats why idk what exactly to think of here

OpenStudy (anonymous):

@aaronq

OpenStudy (aaronq):

you can solve for the molar solubility for all of them, notice the stoichiometry though, for \(CaF_2\), you have Ksp=\(x^3=5.3*10^{-9}\) solubility=x=0.0017435 M check the rest, but this is probably the highest

OpenStudy (somy):

hmmmmm

OpenStudy (aaronq):

i just realized i did that wrong, it's supposed to be \(Ksp=4x^3\)

OpenStudy (anonymous):

But wouldn't it be set as (x)(2x)^2? Which means you would have to divide 5.3 by 4 then take the cube root?

OpenStudy (anonymous):

Yeahhh haha

OpenStudy (aaronq):

haha good eye

OpenStudy (anonymous):

It's the same value as CuCl.

OpenStudy (anonymous):

And CuCl has less ions.

OpenStudy (somy):

i think i should go for A levels on my own i feel so stupid T_T

OpenStudy (aaronq):

hmm, it asks for the cation with the highest solubility... I think \(Ca^{2+}\) then? Try this on paper, choose the same concentration for Cl in CuCl and for F in CaF2. plug it into the Ksp and solve for the cation concentration

OpenStudy (aaronq):

CuCl \(Ksp = 1.2*10^{-6}=[Cu^+][0.1]\) CaF2 \(Ksp = 5.3*10^{-9}=[Ca^{2+}][0.1]^2\) see which is higher

OpenStudy (somy):

0.1 you guys are taking as of anion part?

OpenStudy (aaronq):

yeah

OpenStudy (somy):

interestinggg is it always like that or only when the value for it is not given?>

OpenStudy (anonymous):

Wait why .1??

OpenStudy (aaronq):

well by choosing an arbitrary concentration for the anion we can compare the values, you could've chosen any number

OpenStudy (somy):

lol u just realized?

OpenStudy (somy):

any number?

OpenStudy (aaronq):

well any reasonable number

OpenStudy (aaronq):

you're not gonna use like \(6.022*10^{23}\)

OpenStudy (somy):

so that's as long as i take it as constant value everywhere?

OpenStudy (aaronq):

for examples like this it works

OpenStudy (somy):

noted xD thnx @seehearcreate

OpenStudy (somy):

u get it?

OpenStudy (aaronq):

i just checked it, it still gives Cu^+ as a higher concentration.

OpenStudy (somy):

so that means it'll form ppt

OpenStudy (aaronq):

it means that you can dissolve more Cu^+ than Ca^2+. I'm kinda puzzled, if you say that you chose CuCl and marked it wrong.

OpenStudy (somy):

im lost

OpenStudy (aaronq):

The x for CuCl is actually x=0.0010954 and for CaF2 is x=0.0010983 slightly larger, but larger nonetheless

OpenStudy (aaronq):

i guess within sig figs, they're the same though

OpenStudy (somy):

when i did what u said earlier basically in BaSO4 Ksp = 1.1*10^-10 Ba=1.1*10^-9 CaCO3 Ksp = 2.8*10^-9 Ca= 2.8*10^-8 CaF2 Ksp = 5.3*10^-9 Ca= 5.3*10^-7 CuCl Ksp = 1.2*10^-6 Cu= 1.2*10^-5

OpenStudy (aaronq):

it's more accurate to use the molar solubility, since they vary in stoichiometry. If you chose CuCl and marked it wrong, then it is definitely \(Ca^{2+}\).

OpenStudy (anonymous):

Very sorry! I had to leave to pick up my brother from the bus stop. Yeah, it marked CuCl wrong, I didn't understand it either. I'll try the CaF2.

OpenStudy (somy):

wops

OpenStudy (somy):

wait guys i think i got it

OpenStudy (somy):

this is shown in my book idk if its the thing we need here or not

OpenStudy (anonymous):

OHH. It's a proportion.

OpenStudy (aaronq):

even when you find the solubility, you'd be multiplying them by 1 (thats the coefficient). So nothing much is changed

OpenStudy (anonymous):

Hold on.

OpenStudy (somy):

so did that pic help or it was off topic?

OpenStudy (aaronq):

its definitely on topic, but it doesnt change anything :S

OpenStudy (anonymous):

If anything, that makes the value smaller...

OpenStudy (somy):

hmmmmmm

OpenStudy (anonymous):

I'm about to pound my head against a wall.

OpenStudy (somy):

me2

OpenStudy (somy):

when i did like the book said this is what im getting BaSO4 Ksp = 1.1*10^-10 1.21*10^-20 CaCO3 Ksp = 2.8*10^-9 7.84*10^-18 CaF2 Ksp = 5.3*10^-9 5.62*10^-17 CuCl Ksp = 1.2*10^-6 1.44*10^-12

OpenStudy (anonymous):

Yeah, CuCl still has the highest solubility.

OpenStudy (somy):

why u say that CuCl has highest solubility?

OpenStudy (somy):

O!!!!!!!!!!!!!!!!!! waaait look at another example sec i'll take a pic

OpenStudy (somy):

here

OpenStudy (anonymous):

That isn't changing anything, since you square root Ksp in most of them anyway.

OpenStudy (somy):

ugh

OpenStudy (somy):

i tried this way and the powers in each increase but in CaF2 is stays -9 as it was originally

OpenStudy (anonymous):

@Loser66

OpenStudy (somy):

@Kainui save us T_T

OpenStudy (somy):

BaSO4 Ksp = 1.1*10^-10 Ba+= 1.05*10^-5 CaCO3 Ksp = 2.8*10^-9 Ca+2= 5.29*10^-5 CaF2 Ksp = 5.3*10^-9 Ca+2= 7.49*10^-9 CuCl Ksp = 1.2*10^-6 Cu+= 1.09*10^-3

OpenStudy (somy):

this is what im getting

OpenStudy (anonymous):

I just got feedback, the CaF2 is right. Though how in the heck it is I don't understand. I still want to figure this out.

OpenStudy (somy):

im trying to find connection

OpenStudy (somy):

somehow in all of my calculations answers are in similar trend with BaSO4, CaCO3 and CuCl ONLY CaF2 is totally off trend

OpenStudy (anonymous):

Yep. Yet, CaF2 was correct. I'm wondering if the question was faulty. Someone at corporate falling asleep on their calculator.

OpenStudy (somy):

but to think of this all other answers seem way of more then Ksp only in CaF2 i see only a slight increase

OpenStudy (somy):

i think it makes sense

OpenStudy (somy):

O_O

OpenStudy (somy):

the question is not asking which is 100% soluble right?

OpenStudy (anonymous):

noooo haha. Just which one is most soluble. Which in this case, it is CuCl, no matter what.

OpenStudy (somy):

idk why u think its CuCl but now im kinda sure that its CaF2

OpenStudy (somy):

I mean remember the precipitation question?

OpenStudy (somy):

if value is less then K then no ppt if value more then K then there will be ppt

OpenStudy (somy):

the more the ppt the less soluble

OpenStudy (somy):

in all 3 cases the values are very much bigger then K thus we'd expect more ppt which means less soluble but in CaF2 the value is close to K meaning its even close to being less than K so that means the amount of ppt formed is way less than in the other 3 thus more soluble than the other 3

OpenStudy (somy):

idk if im making sense =_=

OpenStudy (anonymous):

No, larger values of K produce a higher solubility. When k=1, that means the solution is saturated.

OpenStudy (somy):

maybe im trying to connect noncollectable stuff

OpenStudy (somy):

its 1 am night and i can't fall asleep unless this thing is done ~_~

OpenStudy (aaronq):

When you find the molar solubilities, \([Cu^+]\)=0.0010954=0.0011 and \([Ca^{2+}]\)=0.0010983=0.0011, you find that they differ a little bit. But not enough within the given significant digits - so the question is flawed in that sense.

OpenStudy (somy):

i'm off to sleep, mb in the morning i'll get it

OpenStudy (kainui):

I'm pretty sure this is right, but for some reason I'm doubting myself and can't seem to find anywhere that says this on the internet to check myself. At any rate here goes. All of them only have 2 ions they can dissociate into while CaF2 becomes 3. Even though this might seem insignificant compared to CuCl, it actually ends up being what changes everything. CaF2 Ksp = 5.3*10^-9 CuCl Ksp = 1.2*10^-6 Let's calculate the concentration of the ions in solution for CaF2 and CuCl. Actually, we don't have to, you've already done it. [Cu+] = 1.09545*10^-3 [Ca2+]=1.01961*10^-3 Even at this level it appear a close fight, but Copper I ions still win by a marginal amount. But wait a second, what's the total concentration in solution of all the other ions like Chlorine and Fluorine? Looking back at the equilibrium equation, we can see for every 1 copper ion we have 1 chloride. But for every Calcium we have 2 fluoride ions in solution! So let's add them up. 2[Cu+] = 2.19089*10^-3 3[Ca2+] = 3.05884*10^-3 That's your total concentration of dissolved ions in solution. There are more for CaF2 than CuCl. Sorry it took me so long to respond, there was a lot to read through and I am distracted easily lol.

OpenStudy (anonymous):

Well slap me and call me sally....thanks so much(:

OpenStudy (aaronq):

Thats good stuff @Kainui though the question asked about the concentration of cations only :S

OpenStudy (kainui):

yehhhhh well so much for that lol.